Classification of Normal Toric

Varieties over a Valuation Ring of

Rank One

Walter Gubler and Alejandro Soto

Received: March 13, 2013 Revised: July 22, 2014

Communicated by Gavril Farkas

Abstract. Normal toric varieties over a field or a discrete valuation ring are classified by rational polyhedral fans. We generalize this classification to normal toric varieties over an arbitrary valuation ring of rank one. The proof is based on a generalization of Sumihiro’s theorem to this non-noetherian setting. These toric varieties play an important role for tropicalizations.

2010 Mathematics Subject Classification: 14M25, 14L30, 13F30

Contents

1 Introduction  172
2 Divisors on varieties over the valuation ring  175
3 Toric schemes over valuation rings  181
4 The cone of a normal affine toric variety  183
5 Construction of the Cartier divisor  187
6 Linearization and immersion into projective space  191
7 Proof of Sumihiro’s theorem  194

1 Introduction

Toric varieties over a field have been studied since the 70’s. Their geometry is completely determined by the convex geometry of rational polyhedral cones. This gives toric geometry an important role in algebraic geometry for testing conjectures. There are many good references for them, for instance Cox–Little– Schenk [6], Ewald [7], Fulton [8], Kempf–Knudsen–Mumford–Saint-Donat [15] and Oda [21]. Although in these books, toric varieties are defined over an algebraically closed field, the main results hold over any field. Motivated by compactification and degeneration problems, Mumford considered in [15, Chapter IV] normal toric varieties over discrete valuation rings. A similar motivation was behind Smirnov’s paper [24] on projective toric varieties over discrete valuation rings. In the paper of Burgos–Philippon–Sombra [5], toric varieties over discrete valuation rings were considered for applications to an arithmetic version of the famous Bernstein–Kusnirenko–Khovanskii theorem in toric geometry. The restriction to discrete valuation rings is mainly caused by the use of standard methods from algebraic geometry requiring noetherian schemes. At the beginning of the new century, tropical geometry emerged as a new branch of mathematics (see [18] or [19]). We fix now a valued field (K, v) with value group Γ := v(K^{×}) ⊂ ℝ. The Bieri–Groves theorem shows that the tropicalization of a closed d-dimensional subvariety of a split torus T := (𝔾_{m}^{n})_{K} over K is a finite union of Γ-rational d-dimensional polyhedra in ℝ^{n}. Moreover, this tropical variety is the support of a weighted polyhedral complex of pure dimension d such that the canonical tropical weights satisfy a balancing condition in every face of codimension 1. The study of the tropical weights leads naturally to toric schemes over the valuation ring K^{∘} of K (see [14] for details). A 𝕋-toric scheme 𝒴 over K^{∘} is an integral separated flat scheme over K^{∘} containing T as a dense open subset such that the translation action of T on T extends to an algebraic action of the split torus 𝕋 := (𝔾_{m}^{n})_{K^{∘}} on 𝒴 . If a 𝕋-toric scheme is of finite type over K^{∘}, then we call it a 𝕋-toric variety. There is an affine 𝕋-toric scheme 𝒱_{σ} associated to any Γ-admissible cone σ in ℝ^{n} × ℝ_{+}. This construction is similar to the classical theory of toric varieties over a field, where every rational polyhedral cone in ℝ^{n} containing no lines gives rise to an affine T -toric variety. The additional factor ℝ_{+} takes the valuation v into account and Γ-admissible cones are cones containing no lines satisfying a certain rationality condition closely related to the Γ-rationality in the Bieri– Groves theorem. If Σ is a fan in ℝ^{n} × ℝ_{+} of Γ-admissible cones, then we call Σ a Γ-admissible fan. By using a gluing process along common subfaces, we get an associated 𝕋-toric scheme 𝒴_{Σ} over K^{∘} with the open affine covering (𝒱_{σ})_{σ∈Σ}. We refer to §3 for precise definitions. These normal 𝕋-toric schemes 𝒴_{Σ} are studied in [14]. Many of the properties of toric varieties over fields hold also for 𝒴_{Σ}. Rohrer considered in [22] toric schemes X_{Π} over an arbitrary base S associated to a rational fan Π in ℝ^{n} containing no lines. If we restrict to the case S = Spec(K^{∘}), then Rohrer’s toric schemes are a special case of the above 𝕋-toric schemes as we have X_{Π} = 𝒴_{Π×ℝ_{+}}. Note that the cones of Π×ℝ_{+} are preimages of cones in ℝ^{n} with respect to the canonical projection ℝ^{n} × ℝ_{+} → ℝ^{n} and hence the fans Π×ℝ_{+} form a very special subset of the set of Γ-admissible fans in ℝ^{n} × ℝ_{+}. As a consequence, Rohrer’s toric scheme X_{Π} is always obtained by base change from a corresponding toric scheme over Spec(ℤ) while this is in general not true for 𝒴_{Σ}. The generic fibre of 𝒴_{Σ} is the toric variety over K associated to the fan formed by the recession cones of all σ ∈ Σ, but the special fibre of 𝒴_{Σ} need not be a toric variety. In fact, the special fibre of 𝒴_{Σ} is a union of toric varieties corresponding to the vertices of the polyhedral complex Σ ∩ (ℝ^{n} × {1}). On the other hand, every fibre of the toric scheme X_{Π} is a toric variety associated to the same fan Π. This leads to the natural question if every normal 𝕋-toric variety 𝒴 over the valuation ring K^{∘} is isomorphic to 𝒴_{Σ} for a suitable Γ-admissible fan Σ in ℝ^{n} × ℝ_{+}. This classification is possible in the classical theory of normal toric varieties over a field and also in the case of normal toric varieties over a discrete valuation ring. First, one shows that every affine normal 𝕋-toric variety over a field or a discrete valuation ring is of the form 𝒱_{σ} for a rational cone σ in ℝ^{n} × ℝ_{+} containing no lines and then one uses Sumihiro’s theorem which shows that every point in 𝒴 has a 𝕋-invariant affine open neighbourhood (see [15]). Sumihiro proved his theorem over a field in [25]. In [15, Chapter IV], the arguments were extended to the case of a discrete valuation ring. The proof of Sumihiro’s theorem relies on noetherian techniques from algebraic geometry. Now we describe the structure and the results of the present paper. For the generalization of the above classification to normal 𝕋-toric varieties 𝒴 over an arbitrary valuation ring K^{∘} of rank 1, one needs a theory of divisors on varieties over K^{∘}. This will be done in §2. First, we recall some basic facts about normal varieties over K^{∘} due to Knaf [17]. Then we define the Weil divisor associated to a Cartier divisor D and more generally a proper intersection product of D with cycles. This is based on the corresponding intersection theory of Cartier divisors on admissible formal schemes over K^{∘} given in [12]. In §3, we recall the necessary facts for toric varieties over K^{∘} which were proved in [14]. In §4, we show the following classification for affine normal toric varieties:

Theorem 1. If v is not a discrete valuation, then the map σ ↦ 𝒱_{σ} defines a bijection between the set of those Γ-admissible cones σ in ℝ^{n} ×ℝ_{+} for which the vertices of σ ∩ (ℝ^{n} × {1}) are contained in Γ^{n} × {1} and the set of isomorphism classes of normal affine 𝕋-toric varieties over the valuation ring K^{∘}.

Similarly to the classical case, the proof uses finitely generated semigroups in the character lattice of T and duality of convex polyhedral cones. The new ingredient here is an approximation argument concluding the proof of Theorem 1. The additional condition for the vertices of the Γ-admissible cone σ is equivalent to the property that the affine 𝕋-toric scheme 𝒱_{σ} is of finite type over K^{∘} meaning that 𝒱_{σ} is a 𝕋-toric variety over K^{∘} (see Proposition 3.3). If v is a discrete valuation, then 𝒱_{σ} is always a 𝕋-toric variety over K^{∘} and hence the condition on the vertices has to be omitted to get the bijective correspondence in Theorem 1. For the globalization of the classification, the main difficulty is the generalization of Sumihiro’s theorem. The proof follows the same steps working in the case of fields or discrete valuation rings (see [15], proof of Theorem 5 in Chapter I and §4.3). In §5, we show that for every non-empty affine open subset 𝒰_{0} of a normal 𝕋-toric variety 𝒴 over the valuation ring K^{∘} of rank one, the smallest open 𝕋-invariant subset 𝒰 containing 𝒰_{0} has an effective Cartier divisor D with support equal to 𝒰 \ 𝒰_{0}. This is rather tricky in the non-noetherian situation and it is precisely here, where we use the results on divisors from §2. In §6, we use the Cartier divisor D constructed in the previous section to show that 𝒪(D) is an ample invertible sheaf with a 𝕋-linearization. This leads to a 𝕋-equivariant immersion of 𝒰 into a projective space over K^{∘} on which 𝕋-acts linearly. It remains to prove Sumihiro’s theorem for projective 𝕋-toric subvarieties of a projective space over K^{∘} on which 𝕋-acts linearly. This variant of Sumihiro’s theorem will be proved in §7 and relies on properties of such nonnecessarily normal projective 𝕋-toric varieties given in [14, §9]. We get the following generalization of Sumihiro’s theorem:

Theorem 2. Let v be a real valued valuation with valuation ring K^{∘} and let 𝒴 be a normal 𝕋-toric variety over K^{∘}. Then every point of 𝒴 has an affine open 𝕋-invariant neighborhood.

As an immediate consequence, we will obtain our main classification result:

Theorem 3. If v is not a discrete valuation, then the map Σ ↦ 𝒴_{Σ} defines a bijection between the set of fans in ℝ^{n} × ℝ_{+}, whose cones are as in Theorem 1, and the set of isomorphism classes of normal 𝕋-toric varieties over K^{∘}.

If v is a discrete valuation, then we have to omit the additional condition on the vertices of the cones again to get a bijective correspondence in Theorem 3. The authors thank the referee for his comments and suggestions.

Notation

For sets, in A ⊂ B equality is not excluded and A\B denotes the complement of B in A. The set of non-negative numbers in ℤ, ℚ or ℝ is denoted by ℤ_{+}, ℚ_{+} or ℝ_{+}, respectively. All the rings and algebras are commutative with unity. For a ring A, the group of units is denoted by A^{×}. A variety over a field k is an irreducible and reduced scheme which is separated and of finite type over k. See §2 for the definition of varieties over a valuation ring. In the whole paper, we fix a valued field (K, v) which means here that v is a valuation on the field K with value group Γ := v(K^{×}) ⊂ ℝ. Note that K is not required to be algebraically closed or complete and that its valuation can be trivial. We have a valuation ring K^{∘} := {x ∈ K | v(x) ≥ 0} with maximal ideal K^{∘∘} := {x ∈ K | v(x) > 0} and residue field K̃ := K^{∘}/K^{∘∘}. We denote by \overline{K} an algebraic closure of K.

Let M be a free abelian group of rank n with dual N := Hom(M, ℤ). For u ∈ M and ω ∈ N, the natural pairing is denoted by ⟨u, ω⟩ := ω(u) ∈ ℤ. For an abelian group G, the base change is denoted by M_{G} := M ⊗_{ℤ} G; for instance M_{ℝ} = M ⊗_{ℤ} ℝ. The split torus over K^{∘} of rank n with generic fibre T = Spec(K[M]) is given by 𝕋 = Spec(K^{∘}[M]), therefore M can be seen as the character lattice of T and N as the group of one parameter subgroups. For u ∈ M, the corresponding character is denoted by χ^{u}.

2 Divisors on varieties over the valuation ring

The goal of this section is to recall some facts about divisors on varieties over the valuation ring K^{∘} of the valued field (K, v) with value group Γ ⊂ ℝ. The problem here is that K^{∘} need not be noetherian and so we cannot use the usual constructions from algebraic geometry. Instead we will adapt the intersection theory with Cartier divisors on admissible formal schemes from [12] to our algebraic framework. We will start with some topological considerations of varieties over K^{∘}. As we mainly focus on normal varieties in this paper, we will gather some results of Knaf relating the local rings of such varieties to valuation rings. This will be useful to define the Weil divisor associated to a Cartier divisor. If the variety is not normal then we will pass to the formal completion along the special fibre to define the associated Weil divisor according to [12]. This construction will be explained in 2.7–2.9 and is rather technical. In Proposition 2.11, we will show that both constructions agree for normal varieties over K^{∘}. As a consequence, we will obtain a proper intersection theory of Cartier divisors with cycles on a (normal) variety over K^{∘}. The relevant properties are listed in 2.10–2.16. This intersection theory will be used in the proof of the generalization of Sumihiro’s theorem given in §5–7. There we will construct a torus-invariant ample divisor on any normal toric variety over K^{∘} which will be helpful to reduce Sumihiro’s theorem to an easier projective variant.

2.1. A variety over K^{∘} is an integral scheme which is of finite type and separated over K^{∘}. By [14, Lemma 4.2] such a variety 𝒴 is flat over K^{∘}. We have Spec(K^{∘}) = {η, s}, where the generic point η (resp. the special point s) is the zero-ideal (resp. the maximal ideal) in K^{∘}. We get the generic fibre 𝒴_{η} as a variety over K and the special fibre 𝒴_{s} as a separated scheme of finite type over K̃. The variety 𝒴 is called normal if all the local rings 𝒪_{𝒴 ,y} are integrally closed.

Proposition 2.2. A variety 𝒴 over K^{∘} is a noetherian topological space. If d := dim(𝒴_{η}), then every irreducible component of the special fibre has also dimension d. If 𝒴_{s} is non-empty and if v is non-trivial, then the topological dimension of 𝒴 is d + 1. If 𝒴_{s} is empty or if v is trivial, then 𝒴 = 𝒴_{η}.

Proof. The set 𝒴 over K^{∘} is the union of 𝒴_{η} and 𝒴_{s}. This proves the first claim. The second claim follows from flatness of 𝒴 over K^{∘}. The other claims are now obvious.

The following facts about normal varieties over a valuation ring follow from a paper by Knaf [17].

Proposition 2.3. Let 𝒴 be a normal variety over K^{∘}. Then the following properties hold:

(a) For y ∈ 𝒴 , the local ring 𝒪_{𝒴 ,y} is a valuation ring if and only if y is either a dense point of a divisor of the generic fibre 𝒴_{η} or y is a generic point of 𝒴_{s} or 𝒴_{η}.

(b) If y is a dense point of a divisor of the generic fibre, then 𝒪_{𝒴 ,y} is a discrete valuation ring.

(c) If y is a generic point of the special fibre 𝒴_{s}, then 𝒪_{𝒴 ,y} is the valuation ring of a real-valued valuation v_{y} extending v such that Γ is of finite index in the value group of v_{y}.

(d) If 𝒴 = Spec(A), then A = ⋂_{y} 𝒪_{𝒴 ,y} , where y ranges over all points from (b) and (c).

Proof. Since 𝒴_{η} is a normal variety over the field K, it is regular in codimension 1 and hence (b) follows. The claims (a) and (c) follow from [17, Theorem 2.6], where the set of valued points of a normal variety over an arbitrary valuation ring are classified. It remains to prove (d). By [17, Theorem 2.4], the integral domain A is integrally closed and coherent. It follows from [17, 1.3] that A is a Prüfer v-multiplication ring and hence (d) is a consequence of [17, 1.5].

2.4. Let 𝒴 be a variety over K^{∘} with generic fibre Y. A horizontal cycle Z on 𝒴 is a cycle on Y, i.e. Z is a ℤ-linear combination of closed subvarieties W of Y. The support supp(Z) is the union of all closures \overline{W} in 𝒴 , where W ranges over all closed subvarieties with non-zero coefficients. Such W ’s are called prime components of the horizontal cycle Z. If the closure of every prime component of Z in 𝒴 has dimension k (resp. codimension p), then we say that the horizontal cycle Z of 𝒴 has dimension k (resp. codimension p). A vertical cycle V on 𝒴 is a cycle on 𝒴_{s} with real coefficients, i.e. V is an ℝ-linear combination of closed subvarieties W of 𝒴_{s}. The support and prime components are defined as usual. We say that the vertical cycle V of 𝒴 has dimension k (resp. codimension p) if every prime component of V is a closed subvariety of 𝒴_{s} of dimension k (resp. of codimension p in 𝒴 ). A cycle 𝒵 on 𝒴 is a formal sum of a horizontal cycle Z and a vertical cycle V. The support of 𝒵 is supp(𝒵) := supp(Z) ∪ supp(V). If the horizontal part Z and the vertical part V of 𝒵 both have dimension k (resp. codimension p), then we say that 𝒵 has dimension k (resp. codimension p). We say that a cycle is effective if the multiplicities in all its prime components are positive.

2.5. If φ : 𝒴 ′ → 𝒴 is a flat morphism of varieties over K^{∘}, then we define the pull-back φ^{∗}(𝒵) of a cycle 𝒵 on 𝒴 by using flat pull-back of the horizontal and vertical parts. The resulting cycle φ^{∗}(𝒵) of 𝒴 ′ keeps the same codimension as 𝒵. Similarly, we define the push-forward of a cycle with respect to a proper morphism of varieties over K^{∘}. This preserves the dimension of the cycles.

2.6. We recall that the support supp(D) of a Cartier divisor D on 𝒴 is the complement of the union of all open subsets 𝒰 for which D is given by an invertible element in 𝒪(𝒰 ). Clearly, supp(D) is a closed subset of 𝒴 . We say that the Cartier divisor D intersects a cycle 𝒵 of 𝒴 properly, if for every prime component W of 𝒵, we have

codim(supp(D) ∩ \overline{W} , 𝒴 ) ≥ codim(\overline{W} , 𝒴 ) + 1.

2.7. We are going to define the associated Weil divisor cyc(D) of a Cartier divisor D on the variety 𝒴 over K^{∘}. The horizontal part of cyc(D) is the Weil divisor corresponding to the Cartier divisor D|_{Y} on the generic fibre Y of 𝒴 . Thus we just need to construct the vertical part of cyc(D). This will be done by using the corresponding construction for admissible formal schemes over the completion of K^{∘} given in [12]. This is technically rather demanding and we will freely use the terminology and the results from [12]. The reader might skip the details below in a first read trusting that the algebraic intersection theory with Cartier divisors works in the usual way. In fact, we are dealing mostly with normal varieties over K^{∘} in this paper and then one can use Proposition 2.11 to define the multiplicities ord(D, V) of cyc(D) in an irreducible component V of 𝒴_{s} without bothering about admissible formal schemes. To define the vertical part of cyc(D), we may assume that v is non-trivial. Since the special fibre remains the same by base change to the completion of K^{∘}, we may also assume that v is complete. Let 𝒴̂ be the formal completion of 𝒴 along the special fibre (see [26, §6]). This is an admissible formal scheme over K^{∘} with special fibre equal to 𝒴_{s}. We will denote its generic fibre by Y^{∘} which is an analytic subdomain of the analytification Y^{an} of Y. Note that Y^{∘} may be seen as the set of potentially integral points (see [14, 4.9–4.13] for more details). We have a morphism 𝒴̂ → 𝒴 of locally ringed spaces and using pullback, we see that the Cartier divisor D induces a Cartier divisor D̂ on 𝒴̂. By [12, Definition 3.10], we have a Weil divisor cyc(D̂) on 𝒴̂. It follows from [12, Proposition 6.2] that the analytification of the Weil divisor cyc(D|_{Y}) restricts to the horizontal part of cyc(D̂). We define the vertical part of cyc(D) as the vertical part of cyc(D̂). The multiplicity of cyc(D) in an irreducible component V of 𝒴_{s} is denoted by ord(D, V). Since ord(D, V) is linear in D, the map D ↦ cyc(D) is a homomorphism from the group of Cartier divisors to the group of cycles of codimension 1. It follows from the definitions that cyc(D) is an effective cycle if D is an effective Cartier divisor. For convenience of the reader, we recall the definition of ord(D, V) in more details to make these statements obvious.

2.8. First, we assume that K is algebraically closed and that v is complete. We repeat that we (may) assume v non-trivial. To define ord(D, V), we may restrict our attention to an affine neighbourhood of the generic point ζ_{V} of V where D is given by a single rational function f. Hence we may assume 𝒴 = Spec(A) and D = div(f). Then 𝒴̂ is the formal affine spectrum of the ν-adic completion Â of A for any non-zero element ν in the maximal ideal of K^{∘}. We note that 𝒜 := Â ⊗_{K^{∘}} K is a K-affinoid algebra and we have that Y^{∘} is the Berkovich spectrum ℳ (𝒜) of 𝒜. Since Y^{∘} is an analytic subdomain of Y^{an}, we conclude that Y^{∘} is reduced (see [1, Proposition 3.4.3] for a proof). Let 𝒜^{∘} be the K^{∘}-subalgebra of power bounded elements in 𝒜. Then 𝔜′′ := Spf(𝒜^{∘}) is an admissible formal affine scheme over K^{∘} with reduced special fibre and we have a canonical morphism 𝔜′′ → 𝒴̂. The restriction of this morphism to the special fibres is finite and surjective (see [14, 4.13] for the argument). In particular, there is a generic point y′′ of 𝔜′′_{s} over ζ_{V}. It follows from [1, Proposition 2.4.4] that there is a unique ξ′′ in the generic fibre Y^{∘} of 𝔜′′ with reduction y′′. Recall that the elements of Y^{∘} are bounded multiplicative seminorms on 𝒜. Since y′′ is a generic point of the special fibre of 𝔜′′ = Spf(𝒜^{∘}), the seminorm corresponding to ξ′′ is in fact an absolute value with valuation ring equal to 𝒪_{𝔜′′,y′′}. This follows from [3, Proposition 6.2.3/5], where necessary and sufficient conditions are given for the supremum seminorm to be a valuation. We use it to define

ord(D, V) := − ∑_{y′′}[K̃(\overline{y′′}) : K̃(V)] log |f(ξ′′)|,  (1)

where y′′ is ranging over all generic points of 𝔜′′_{s} mapping to the generic point ζ_{V} of V.

2.9. If K is not algebraically closed, then we perform base change to ℂ_{K}. The latter is the completion of the algebraic closure of the completion of K. This is the smallest algebraically closed complete field extending the valued field (K, v), and the residue field ℂ̃_{K} is the algebraic closure of K̃ [3, §3.4.1]. Again, we may assume 𝒴 = Spec(A) and D = div(f) for a rational function f on 𝒴 . Let 𝒴 ′ be the base change of 𝒴 to the valuation ring ℂ_{K}^{∘} of ℂ_{K}. Let (𝒴′_{j})_{j=1,...,r} be the irreducible components of 𝒴 ′. Our goal is to define ord(D, V) in the irreducible component V of 𝒴_{s}. The definition will be determined by the two guidelines that cyc(D) should be invariant under base change to ℂ_{K}^{∘} and that this base change should be linear in the irreducible components 𝒴′_{j}. Since we do not assume that a variety is geometrically reduced, the multiplicity m(Y′_{j}, Y′) of the generic fibre Y′_{j} of 𝒴′_{j} in the generic fibre Y′ of 𝒴 ′ has to be considered. Note also that the absolute Galois group Gal(ℂ̃_{K}/K̃) acts transitively on the irreducible components of the base change Vℂ̃_{K} and hence the multiplicity m(V′, Vℂ̃_{K}) is independent of the choice of an irreducible component V′ of Vℂ̃_{K}. We choose an irreducible component V′ of Vℂ̃_{K}. It is also an irreducible component of 𝒴′_{s} and hence there is an irreducible component 𝒴′_{j} containing the generic point ζ_{V′} of V′. For 𝒴′_{j} = Spec(A′_{j}), we proceed as in 2.8. We get an admissible formal scheme 𝔜′′_{j} = Spf((𝒜′′_{j})^{∘}) over ℂ_{K}^{∘} with reduced generic fibre (Y′_{j})^{∘} = ℳ (𝒜′′_{j}) and reduced special fibre (𝔜′′_{j})_{s} with a surjective finite map onto (𝒴′_{j})_{s}. Hence there is at least one generic point y′′_{j} of (𝔜′′_{j})_{s} mapping to ζ_{V′}. Again, there is a unique point ξ′′_{j} ∈ (Y′_{j})^{∘} with reduction y′′_{j}. Then ξ′′_{j} extends to an absolute value with valuation ring 𝒪_{𝔜′′_{j},y′′_{j}} and (1) leads to the definition

ord(D, V) := −\frac{1}{m(V′, Vℂ̃_{K})} ∑_{j} m(Y′_{j}, Y′) ∑_{y′′_{j}}[ℂ̃_{K}(\overline{y′′_{j}}) : ℂ̃_{K}(V′)] log |f(ξ′′_{j})|,

(2)

where Y′_{j} ranges over the irreducible components of Y′ and y′′_{j} ranges over the generic points of (𝔜′′_{j})_{s} lying over the generic point ζ_{V′} of V′. Using the action of Gal(ℂ̃_{K}/K̃), we see that the definition is independent of the choice of the irreducible component V′ of Vℂ̃_{K}. It follows from compatibility of base change and passing to the formal completion along the special fibre that ∑_{V} ord(D, V)V is indeed the vertical part of cyc(D) as defined in 2.7.

The Weil divisor associated to a Cartier divisor has all the expected properties. The proofs follow from the corresponding properties in [12] or [13]. This is illustrated in the proof of the following projection formula:

Proposition 2.10. Let φ : 𝒴 ′ → 𝒴 be a proper morphism of varieties over K^{∘} and let D be a Cartier divisor on 𝒴 such that supp(D) does not contain φ(𝒴 ′). As usual, we define [𝒴 ′ : 𝒴 ] to be the degree of the extension of the fields of rational functions if this degree is finite and 0 otherwise. Then we have

φ_{∗}(cyc(φ^{∗}(D))) = [𝒴 ′ : 𝒴 ]cyc(D).

Proof. The projection formula holds in the generic fibre [9, Proposition 2.3]. We have an induced proper morphism φ̂ : 𝒴̂ ′ → 𝒴̂ of admissible formal schemes over the completion of K^{∘} and hence the projection formula follows for vertical parts from the projection formula for proper morphisms of admissible formal schemes given in [12, Proposition 4.5] and from the compatibility of pushforward and passing to the analytification given in [12, Proposition 6.3].

Proposition 2.11. Assume that v is non-trivial, let 𝒴 be a normal variety over K^{∘} and let V be an irreducible component of 𝒴_{s}. If the Cartier divisor D on 𝒴 is given by the rational function f in a neighbourhood of the generic point y = ζ_{V} of V, then 𝒪_{𝒴 ,y} is a valuation ring for a unique real-valued valuation v_{y} extending v and we have ord(D, V) = v_{y}(f).

Proof. We may assume that 𝒴 = Spec(A) and D = div(f) for a rational function f on 𝒴 . The first claim follows from Proposition 2.3. In the following, we use the notation and the results from 2.9. We have a generic point y′′_{j} of the special fibre of the admissible formal scheme 𝔜′′_{j} := Spf((𝒜′′_{j})^{∘}) lying over y. We have seen in 2.9 that the unique point ξ′′ of the generic fibre of 𝔜′′_{j} mapping to y′′_{j} extends to an absolute value with valuation ring 𝒪_{𝔜′′_{j},y′′_{j}}. We conclude that the valuation ring 𝒪_{𝔜′′_{j},y′′_{j}} dominates the valuation ring 𝒪_{𝒴 ,y} .

Since valuation rings are maximal with respect to dominance of local rings in a given field, we conclude that − log |f(ξ′′)| = v_{y}(f) and hence (2) simplifies to

ord(D, V) := \frac{v_{y}(f)}{m(V′, Vℂ̃_{K})} ∑_{j} m(Y′_{j}, Y′) ∑_{y′′_{j}}[ℂ̃_{K}(\overline{y′′_{j}}) : ℂ̃_{K}(V′)],  (3)

where Y′_{j} ranges over the irreducible components of Y′ and y′′_{j} ranges over the generic points of (𝔜′′_{j})_{s} lying over the generic point ζ_{V′} of V′. By [13, Lemma 4.5], we have

m(V′, Vℂ̃_{K}) = ∑_{j} m(Y′_{j}, Y′) ∑_{y′′_{j}}[ℂ̃_{K}(\overline{y′′_{j}}) : ℂ̃_{K}(V′)]

proving the claim.

Corollary 2.12. The following properties hold for a Cartier divisor D on a normal variety 𝒴 over K^{∘}.

(a) supp(D) = supp(cyc(D)).

(b) The Cartier divisor D is effective if and only if cyc(D) is an effective cycle.

(c) The map D ↦ cyc(D) is an injective homomorphism from the group of Cartier divisors on 𝒴 to the group of cycles of codimension 1 on 𝒴 .

Proof. It follows easily from the definitions that supp(cyc(D)) ⊂ supp(D) and that the Weil divisor associated to an effective Cartier divisor is an effective cycle without assuming normality. If v is trivial, the claims are classical results for divisors on normal varieties over K and so we may assume that v is nontrivial. Then (b) follows from Propositions 2.11 and 2.3. To prove (a), the above shows that by passing to the open subset 𝒴 \ supp(cyc(D)), we may assume that cyc(D) = 0 and hence (a) follows from (b). Similarly, (c) is a consequence of (b).

2.13. The construction of the Weil divisor associated to a Cartier divisor allows us to define a proper intersection product of a Cartier divisor with a cycle. Indeed, let D be a Cartier divisor intersecting the cycle 𝒵 on 𝒴 properly. Then we define the proper intersection product D.𝒵 as a cycle on 𝒴 in the following way: By linearity, we may assume that 𝒵 is a prime cycle W. If W is vertical, then D restricts to a Cartier divisor on W and we define D.W := cyc(D|_{W}) using algebraic intersection theory on the variety W. If W is horizontal, then D restricts to a Cartier divisor on the closure of W in 𝒴 and we define D.W as the associated Weil divisor. Obviously, this proper intersection product is bilinear.

Proposition 2.14. Let D and E be properly intersecting Cartier divisors on 𝒴 which means codim(supp(D)∩supp(E), 𝒴 ) ≥ 2. Then we have D.cyc(E) = E.cyc(D).

Proof. For the horizontal parts, this follows from algebraic intersection theory. For the vertical parts, this follows from the corresponding statement for Cartier divisors on admissible formal schemes given in [12, Theorem 5.9]. Recall that the special fibre doesn’t change after passing to the formal completion of 𝒴 along 𝒴_{s}.

Proposition 2.15. Let φ : 𝒴 ′ → 𝒴 be a flat morphism of varieties over K^{∘} and let D be a Cartier divisor on 𝒴 . Then we have φ^{∗}(cyc(D)) = cyc(φ^{∗}(D)).

Proof. Since φ is flat, the pull-back of D is well-defined as a Cartier divisor and the claim follows from [13, Proposition 4.4(d)], where it is proved for Cartier divisors on admissible formal schemes.

2.16. We say that two cycles 𝒟_{1} and 𝒟_{2} of codimension 1 on the variety 𝒴 over K^{∘} are rationally equivalent if there is a non-zero rational function f on 𝒴 such that 𝒟_{1} − 𝒟_{2} = cyc(div(f)). The first Chow group CH^{1}(𝒴 ) of 𝒴 is defined as the group of cycles of codimension 1 modulo rational equivalence. It follows from Proposition 2.15 that rational equivalence is compatible with flat pull-back. Two Cartier divisors D_{1} and D_{2} on 𝒴 are said to be linearly equivalent if there is a non-zero rational function f on 𝒴 such that D_{1} − D_{2} = div(f). The group of Cartier divisors modulo linear equivalence is isomorphic to Pic(𝒴 ) using the map D ↦ 𝒪(D). We may use rational equivalence to define a refined intersection theory with pseudo divisors on a variety 𝒴 over K^{∘} with the same properties as in [9, Chapter 2]. The proofs follow directly from [12] and [13, §4]. This will not be used in the sequel and so we leave the details to the interested reader.

3 Toric schemes over valuation rings

In this section, (K, v) is a valued field with valuation ring K^{∘} , residue field K̃ and value group Γ := v(K^{×}) ⊂ ℝ. As usual, 𝕋 = Spec(K^{∘}[M]) is the split torus of rank n with generic fibre T and N is the dual of the free abelian group M. We will start with the definition of a 𝕋-toric scheme over K^{∘}. Then we will recall from [14] the basic construction of a normal 𝕋-toric scheme associated to a Γ-admissible fan. The overall goal of this paper is to show that every normal toric variety over K^{∘} arises in this way. At the end of this section, we will study projective toric varieties over K^{∘} with a linear action of the torus. They can also be understood purely in combinatorial terms, but they are not necessarily normal. We will use them in §7 to finish the proof of Sumihiro’s theorem.

Definition 3.1. A 𝕋-toric scheme over the valuation ring K^{∘} is an integral separated flat scheme 𝒴 over K^{∘} such that the generic fibre 𝒴_{η} contains T as an open subset and such that the translation action T ×_{K} T → T extends to an algebraic action 𝕋 ×_{K^{∘}} 𝒴 → 𝒴 over K^{∘}.

A homomorphism (resp. isomorphism) of 𝕋-toric schemes is an equivariant morphism (resp. isomorphism) which restricts to the identity on T. A 𝕋-toric scheme of finite type over K^{∘} is called a 𝕋-toric variety.

Note that if 𝒴 is a 𝕋-toric variety over K^{∘}, then 𝒴_{η} is a T -toric variety over K. In order to construct examples of 𝕋-toric schemes and to see how they can be described by the combinatorics of some objects in convex geometry, we need to introduce and to study the following algebras associated to Γ-admissible cones.

3.2. A cone σ ⊂ N_{ℝ} × ℝ_{+} is called Γ-admissible if it can be written as

σ = ⋂_{i=1}^{k} {(ω, s) ∈ N_{ℝ} × ℝ_{+} | ⟨u_{i}, ω⟩ + sc_{i} ≥ 0} , u_{1}, …, u_{k} ∈ M, c_{1}, …, c_{k} ∈ Γ,

and does not contain a line. For such a cone σ, we define

K[M]^{σ} := { ∑_{u∈M} α_{u}χ^{u} ∈ K[M] | cv(α_{u}) + ⟨u, ω⟩ ≥ 0 ∀(ω, c) ∈ σ}

and 𝒱_{σ} := Spec(K[M]^{σ}). It is easy to see that K[M]^{σ} is an M-graded K^{∘}- algebra and hence we have a canonical 𝕋-action on 𝒱_{σ}.

Proposition 3.3. Let σ be a Γ-admissible cone in N_{ℝ} × ℝ_{+}. Then 𝒱_{σ} is a normal 𝕋-toric scheme over K^{∘}. If v is a discrete valuation, then 𝒱_{σ} is always a 𝕋-toric variety. If v is not a discrete valuation, then 𝒱_{σ} is a 𝕋-toric variety over K^{∘} if and only if the vertices of σ ∩ (N_{ℝ} × {1}) are contained in N_{Γ} × {1}.

Proof. Normality is proven in [14, Proposition 6.10]. If the valuation is discrete, then [14, Proposition 6.7] shows that 𝒱_{σ} is a 𝕋-toric variety. If v is not discrete, then the equivalence in the last claim is proved in [14, Propositions 6.9].

3.4. A Γ-admissible fan Σ in N_{ℝ} × ℝ_{+} is a fan consisting of Γ-admissible cones. Given a Γ-admissible fan Σ, we glue the normal affine 𝕋-toric schemes 𝒱_{σ}, σ ∈ Σ, along the open subschemes coming from their common faces. The result is a normal 𝕋-toric scheme 𝒴_{Σ}. Similarly to the classical case of toric varieties over a field, the properties of the 𝕋-toric schemes 𝒴_{Σ} may be described by the combinatorics of the cones Σ. For details, we refer to [14].

Now we review the construction of projective 𝕋-toric schemes which are not necessarily normal (see [14, §9 ] for more details). These are not all the possible projective toric schemes over K^{∘} but just those which have a linear action of the torus, see [14, Proposition 9.8]. For the corresponding projective toric varieties over a field, we refer to Cox–Little–Schenk [6, §2.1, §3.A] and Gelfand– Kapranov–Zelevinsky [10, Chapter 5].

3.5. Given R ∈ ℤ_{+}, we choose projective coordinates on the projective space ℙ_{K^{∘}}^{R}. Let A = (u_{0}, …, u_{R}) ∈ M^{R+1} and y = (y_{0} : ⋯: y_{R}) ∈ ℙ^{R}(K). The height function of y is defined as

a : {0, ..., R} → Γ ∪ {∞}, j ↦ a(j) := v(y_{j}).

The action of 𝕋 on ℙ_{K^{∘}}^{R} is given by

(t, x) ↦ (χ^{u_{0}}(t)x_{0} : ⋯: χ^{u_{R}}(t)x_{R}).

We define the projective toric variety 𝒴_{A,a} to be the closure of the orbit T y. The generic fibre Y_{A,a} is a toric variety respect to the torus T/Stab(y). It follows from [14, 9.2] that 𝒴_{A,a} is a 𝕋-toric variety over K^{∘} with respect to the split torus over K^{∘} with generic fibre T/Stab(y). The weight polytope Wt(y) is defined as the convex hull of A(y) := {u_{j} | a(j) < ∞}. The weight subdivision Wt(y, a) is the polytopal complex with support Wt(y) obtained by projecting the faces of the convex hull of {(u_{j}, λ_{j}) ∈ M_{ℝ} × ℝ_{+} | j = 0, …, R; λ_{j} ≥ a(j)}. We will see in the next result that the orbits of 𝒴_{A,a} can be read off from the weight subdivision.

Proposition 3.6. There is a bijective order preserving correspondence between faces Q of the weight subdivision Wt(y, a) and 𝕋-orbits Z of the special fibre of 𝒴_{A,a} given by

Z = {x ∈ (𝒴_{A,a})_{s} | x_{j} ≠ 0 ⇔ u_{j} ∈ A(y) ∩ Q}.

Proof. This is the content of [14, Proposition 9.12].

4 The cone of a normal affine toric variety

We recall that (K, v) is a valued field with valuation ring K^{∘}, residue field K̃ and value group Γ ⊂ ℝ. Let 𝕋 = Spec(K^{∘}[M]) be the split torus over K^{∘} with generic fibre T. The free abelian group M of rank n is isomorphic to the character group of T. For an element u ∈ M, the corresponding character is denoted by χ^{u}. Let N = Hom(M, ℤ) be the dual abelian group of M. As we have seen in the previous section, a Γ-admissible cone σ in N_{ℝ} × ℝ_{+} induces a normal affine 𝕋-toric scheme 𝒱_{σ} = Spec(K[M]^{σ}). This is a 𝕋-toric variety if and only if the vertices of σ ∩ (N_{ℝ} × {1}) are contained in N_{Γ} × {1} or if v is discrete. In this section, we will show that every normal affine 𝕋-toric variety 𝒴 = Spec(A) has this form proving Theorem 1. We may assume that the valuation is non-trivial as in the classical case of normal toric varieties over a field, the statement is well known (see [15, ch. I, Theorem 1]). The 𝕋-action induces an M-grading A = ⨁_{m∈M} A_{m} on the K^{∘}-algebra A. Since T is an open dense orbit of 𝒴 , we may and will assume that A is a subalgebra of the quotient field K(M) of K[M]. The proof of Theorem 1 will be done in three steps along the classical proof dealing with the additional complications of a non-trivial value group and of a non-noetherian ring A. In Lemma 4.1, we will associate to the M-graded subalgebra A a semigroup S in M × Γ. Then in a second step, we will show in Lemmata 4.2 and 4.3 that S gives rise to a Γ-admissible cone σ in N × ℝ_{+}. Finally, we will prove in Proposition 4.4 that 𝒴 = 𝒱_{σ} by using an approximation argument.

Lemma 4.1. The set S := {(m, v(a)) ∈ M × Γ | aχ^{m} ∈ A\{0}} is a saturated semigroup in M × Γ.

Proof. Obviously, the set S is a semigroup. Let k(m, v(a)) ∈ S for m ∈ M, a ∈ K \ {0} and k ∈ ℤ_{+} \ {0}, i.e. (aχ^{m})^{k} ∈ A. By normality of A, we get aχ^{m} ∈ A and hence (m, v(a)) ∈ S.

Lemma 4.2. There are M-homogeneous generators a_{1}χ^{m_{1}}, …, a_{k}χ^{m_{k}} of A. Moreover, the semigroup S from Lemma 4.1 and the set {(0, 1), (m_{i}, v(a_{i})) | i = 1, …, k} generate the same cone in M_{ℝ} × ℝ.

Proof. Since 𝒴 is a variety over K^{∘}, it is clear that A is a finitely generated K^{∘}-algebra. Using that A is an M-graded algebra, we find generators a_{1}χ^{m_{1}}, …, a_{k}χ^{m_{k}} of A. Obviously, every (m_{i}, v(a_{i})) is contained in the cone generated by S which we denote by cone(S). Since the valuation v is non-trivial, it is clear that (0, 1) ∈ cone(S). It remains to show that the cone generated by S is contained in the cone generated by {(0, 1), (m_{i}, v(a_{i})) | i = 1, …, k}. An element of cone(S) is a finite sum ∑_{j} α_{j}(u_{j}, v(b_{j})) with α_{j} ∈ ℝ_{+} and (u_{j}, v(b_{j})) ∈ S with b_{j}χ^{u_{j}} ∈ A. Using the above generators, we get

b_{j}χ^{u_{j}} = λ^{(j)}(a_{1}χ^{m_{1}})^{l_{1}^{(j)}} ⋯ (a_{k}χ^{m_{k}})^{l_{k}^{(j)}} for λ^{(j)} ∈ K^{∘} \ {0}, l_{1}^{(j)}, …, l_{k}^{(j)} ∈ ℤ_{+}.

This implies

v(b_{j}) = v(λ^{(j)}) + ∑_{i=1}^{k} l_{i}^{(j)}v(a_{i})

u_{j} = ∑_{i=1}^{k} l_{i}^{(j)}m_{i}.

We conclude that the element ∑_{j} α_{j}(u_{j}, v(b_{j})) of cone(S) is equal to

∑_{j} α_{j} (∑_{i=1}^{k} l_{i}^{(j)}m_{i}, v(λ^{(j)}) + ∑_{i=1}^{k} l_{i}^{(j)}v(a_{i}))

= ∑_{j} α_{j}(0, v(λ^{(j)})) + ∑_{j} ∑_{i} α_{j}l_{i}^{(j)}(m_{i}, v(a_{i}))

= (0, λ) + ∑_{i} λ_{i}(m_{i}, v(a_{i}))

with λ := ∑_{j} α_{j}v(λ^{(j)}) ∈ ℝ_{+} and λ_{i} := ∑_{j} α_{j}l_{i}^{(j)} ∈ ℝ_{+}. This proves the lemma.

Lemma 4.3. The set σ := {(ω, s) ∈ N_{ℝ} × ℝ | ⟨u, ω⟩ + ts ≥ 0 ∀(u, t) ∈ S} is a Γ-admissible cone in N_{ℝ} × ℝ_{+}.

Proof. By definition, σ is the dual cone of the cone generated by S. From Lemma 4.2, we have

σ = ⋂_{i=1}^{k} {(ω, s) ∈ N_{ℝ} × ℝ_{+} | ⟨m_{i}, ω⟩ + sv(a_{i}) ≥ 0} .

It remains to show that σ doesn’t contain a line. Suppose σ contains a line. Then we have ℝ · (ω, t) ⊂ σ for some (ω, t) ∈ N_{ℝ} × ℝ_{+}. Since σ ⊂ N_{ℝ} × ℝ_{+}, we must have t = 0. Therefore the line is of the form ℝ · (ω, 0) ⊂ N_{ℝ} × {0}. For any aχ^{u} ∈ A \ {0}, we have (u, v(a)) ∈ S and hence

0 ≤ ⟨(u, v(a)), (λω, 0)⟩ = λ⟨u, ω⟩ ∀λ ∈ ℝ.

This proves u ∈ ω^{⊥}. Choosing a basis {u_{1}, …, u_{n}} for M such that u_{1}, …, u_{n−1} ∈ ω^{⊥}, we get A ⊂ K[χ^{±u_{1}}, …, χ^{±u_{n−1}}]. On the other hand, 𝒴 is a 𝕋-toric variety and hence the quotient field of A is K(χ^{±u_{1}}, …, χ^{±u_{n}}). This is a contradiction and hence σ doesn’t contain any line. We conclude that σ is Γ-admissible.

Proposition 4.4. Let 𝒴 = Spec(A) be an affine normal 𝕋-toric variety over K^{∘}. Then 𝒴 = 𝒱_{σ} for the Γ-admissible cone σ defined in Lemma 4.3.

Proof. We have to show K[M]^{σ} = A. Take any aχ^{m} ∈ A \ {0}. Since (m, v(a)) ∈ S, we get ⟨m, ω⟩ + t · v(a) ≥ 0 for all (ω, t) ∈ σ and hence aχ^{m} ∈ K[M]^{σ}. This proves A ⊂ K[M]^{σ}. To prove the reverse inclusion, we take aχ^{m} ∈ K[M]^{σ} \ {0}. By definition, (m, v(a)) is in the dual cone σ̌ of σ. Using biduality of convex polyhedral cones (see [8, §1.2]), we conclude that (m, v(a)) is contained in the cone in M_{ℝ} × ℝ generated by S. By Lemma 4.2, we get

(m, v(a)) = κ(0, 1) + ∑_{i=1}^{k} λ_{i}(m_{i}, v(a_{i})), κ, λ_{i} ∈ ℝ_{+}.

From this, we deduce the following equivalent system of equations

m = ∑_{i} λ_{i}m_{i}  (4)
v(a) = κ + ∑_{i} λ_{i}v(a_{i}).  (5)

Now we show that it is always possible to choose all λ_{i} ∈ ℚ_{+}. We may assume that λ_{i} > 0 for all i, otherwise we omit these coefficients. Let b_{1}, …, b_{s} be a basis in ℚ^{k} for the solutions of the homogeneous equation associated to (4) and let μ ∈ ℚ^{k} be a particular solution for (4). We will use the coordinates (b_{j}^{(1)}, …, b_{j}^{(k)}) for the vector b_{j} of the basis. The space of solutions 𝕃 is given by

𝕃 = {μ + ∑_{j=1}^{s} ρ_{j}b_{j} | ρ_{j} ∈ ℝ, j = 1, …, s}.

Since λ := (λ_{i}) ∈ ℝ_{+}^{k} is a solution of (4), there exist ρ_{j} ∈ ℝ (j = 1, …, s) such that

λ = μ + ∑_{j} ρ_{j}b_{j} .

Now choose ρ̂_{j} ∈ ℚ close to ρ_{j}, i.e.

ρ_{j} = ρ̂_{j} + ϵ_{j}

with |ϵ_{j}| small. Then

λ̂ = μ + ∑_{j} ρ̂_{j}b_{j}

is also a solution of (4) in ℚ^{k} which is close to λ. In particular, we may choose |ε_{j}| so small that all λ̂_{i} > 0. Explicitly we have

( λ_{1} ; ⋮ ; λ_{k} ) = ( λ̂_{1} ; ⋮ ; λ̂_{k} ) + ∑_{j=1}^{s} ϵ_{j}b_{j}.

Inserting this in (5), we get

v(a) = κ + ∑_{i} (λ̂_{i} + ∑_{j} ϵ_{j}b_{j}^{(i)}) v(a_{i})

= κ + ∑_{i} λ̂_{i}v(a_{i}) + ∑_{i} (∑_{j} ϵ_{j}b_{j}^{(i)}) v(a_{i}).

With α := ∑_{i} (∑_{j} ϵ_{j}b_{j}^{(i)}) v(a_{i}) = ∑_{j} ϵ_{j} ∑_{i} b_{j}^{(i)}v(a_{i}), we get

v(a) = κ + α + ∑_{i} λ̂_{i}v(a_{i}).

It is easy to see that we may choose ϵ_{1}, …, ϵ_{s} in a small neighbourhood of 0 such that κ + α ≥ 0. We conclude that it is possible to choose the coefficients in (4) rational and we have

(m, v(a)) = (κ + α)(0, 1) + ∑_{i=1}^{k} λ̂_{i}(m_{i}, v(a_{i})), κ + α ∈ ℝ_{+}, λ̂_{i} ∈ ℚ_{+}. (6)

The above shows that (m, v(a)) = (0, κ) + ∑_{i} λ_{i}(m_{i}, v(a_{i})) with λ_{i} ∈ ℚ_{+} and κ ∈ ℝ_{+}. Let R be a positive integer such that Rλ_{i} ∈ ℤ_{+} for i = 1, …, k. Then we get

R(m, v(a)) = R(0, κ) + ∑_{i} Rλ_{i}(m_{i}, v(a_{i})).

This proves in particular that Rκ ∈ Γ. Since (0, Rκ), (m_{i}, v(a_{i})) ∈ S (i = 1, …, k) and Rλ_{i} ∈ ℤ_{+}, we conclude that (Rm, Rv(a)) is also in the semigroup S. This means that there is a b ∈ K with v(b) = v(a^{R}) such that bχ^{Rm} ∈ A. Since a^{R} = ub for a unit in K^{∘}, we get (aχ^{m})^{R} ∈ A. By normality of A, this implies that aχ^{m} ∈ A. We conclude that K[M]^{σ} = A and hence 𝒴 = 𝒱_{σ}.

Proof of Theorem 1. We assume that v is not a discrete valuation. We have seen in Propositions 4.4 and 3.3 that the map σ ↦ 𝒱_{σ} from the set of those Γ-admissible cones in N_{ℝ} × ℝ_{+} for which the vertices of σ ∩ (N_{ℝ} × {1}) are contained in N_{Γ} ×{1} to the set of isomorphism classes of affine normal 𝕋-toric varieties over K^{∘} is surjective. By [14, Proposition 6.24], we can reconstruct the cone σ from the 𝕋-toric scheme 𝒱_{σ} by applying the tropicalization map to the set of integral points of T ∩ 𝒱_{σ} and hence the correspondence is indeed bijective. If v is a discrete valuation, then the same argument works if we omit the additional condition on the vertices of the cones.

5 Construction of the Cartier divisor

Let (K, v) be a valued field with valuation ring K^{∘}, value group Γ = v(K^{×}) ⊂ ℝ and residue field K̃. Let 𝕋 be the split torus of rank n over K^{∘}. In this section, we consider a non-empty affine open subset 𝒰_{0} in a normal 𝕋-toric variety 𝒴 over K^{∘}. The main goal is to construct an effective Cartier divisor D on the smallest 𝕋-invariant open subset 𝒰 of 𝒴 containing 𝒰_{0} with support equal to 𝒰 \ 𝒰_{0}. This will be achieved in Proposition 5.12. We will start by noting in Proposition 5.1 that normality of 𝒴 yields that the complement of 𝒰_{0} is a Weil divisor. Using the divisorial intersection theory from §2, we will deduce in Proposition 5.9 that all translates of this Weil divisor are rationally equivalent on 𝒰 . This will be enough to deduce that the Weil divisor is actually a Cartier divisor on 𝒰 proving Proposition 5.12. In fact, it will also follow in Corollary 5.14 that 𝒪(D) is generated by global sections. This will be important in the next section, where we will show that D is ample and gives rise to an equivariant immersion of 𝒰 to a projective space with a linear 𝕋-action. This will allow us in §7 to reduce the proof of Sumihiro’s theorem to an easier projective variant.

Proposition 5.1. Let 𝒰_{0} be a non-empty affine open subset of a normal variety 𝒴 over K^{∘}. Then every irreducible component of 𝒴 \ 𝒰_{0} has codimension 1 in 𝒴 .

Proof. By removing the irreducible components of 𝒴 \ 𝒰_{0} of codimension 1, we may assume that 𝒴 \ 𝒰_{0} has no irreducible components of codimension 1. Then we have to prove 𝒰_{0} = 𝒴 . We may assume that 𝒴 is an affine variety Spec(A). Using Proposition 2.3(d), we get 𝒪(𝒰_{0}) = 𝒪(𝒴 ) and hence the affine varieties 𝒰_{0} and 𝒴 are equal.

In the following result, we will use the notions introduced in Section 2.

Proposition 5.2. Let p_{2} be the canonical projection of 𝕋×_{K^{∘}} 𝒴 onto the variety 𝒴 over K^{∘} and let 𝒟 be a cycle of codimension 1 in 𝕋 ×_{K^{∘}} 𝒴 . Then there is a cycle 𝒟′ on 𝒴 of codimension 1 such that p_{2}^{∗}(𝒟′) is rationally equivalent to 𝒟.

Proof. We note that irreducible components of (𝕋×_{K^{∘}}𝒴 )_{s} are given by 𝕋_{s}×K̃V with V ranging over the irreducible components of 𝒴_{s}. We conclude that every vertical cycle of codimension 1 in 𝕋 ×_{K^{∘}} 𝒴 is the pull-back of a vertical cycle of codimension 1 in 𝒴 . This reduces the claim to the horizontal parts where it is a standard fact from algebraic intersection theory on varieties over a field [9, Proposition 1.9].

5.3. The generic fibre of 𝕋 is denoted by T. Let T^{∘} be the affinoid torus in T^{an} given by {x ∈ T^{an} | |x_{1}(x)| = ⋯ = |x_{n}(x)| = 1} in terms of torus coordinates x_{1}, …, x_{n} and let 𝒴 be a variety over K^{∘} with generic fibre Y. For t ∈ T^{∘}(K), the reduction t̃ ∈ 𝕋_{s}(K̃) is well-defined. Let i_{t} : 𝒴 → 𝕋 ×_{K^{∘}} 𝒴 be the embedding over 𝒴 induced by the integral point of 𝕋 corresponding to t. We are going to define the pull-back i_{t}^{∗}(𝒵) for every cycle 𝒵 on 𝕋 ×_{K^{∘}} 𝒴 which satisfies the following flatness condition: We assume that every component of the horizontal (resp. vertical) part of 𝒵 is flat over T (resp. 𝕋_{s}). Since i_{t} induces a regular embedding of Y into T ×_{K} Y (resp. of 𝒴_{s} into 𝕋_{s} ×K̃ 𝒴_{s}), the pull-back of the horizontal part (resp. vertical part) of 𝒵 is a well-defined cycle on Y (resp. 𝒴_{s}) (see [9, Chapter 6]). We define i_{t}^{∗}(𝒵) as the sum of these two pull-backs. Clearly, this pull-back keeps the codimension and is linear in 𝒵.

5.4. For t ∈ T^{∘}(K) with coordinates t_{1} := x_{1}(t), …, t_{n} := x_{n}(t), let D_{t_{j}} be the Cartier divisor on 𝕋 ×_{K^{∘}} 𝒴 given by pull-back of div(x_{j} − t_{j}) with respect to the canonical projection onto 𝕋 = (𝔾_{m}^{n})_{K^{∘}}. Let 𝒵 be a cycle on 𝕋 ×_{K^{∘}} 𝒴 satisfying the flatness condition from 5.3. Then we may use the proper intersection product with Cartier divisors from 2.13 to get

(i_{t})_{∗}(i_{t}^{∗}(𝒵)) = D_{t_{1}} … D_{t_{n}}.𝒵 .

Indeed, the flatness condition ensures that the right hand side is a proper intersection product and hence the claim follows from [9, Example 6.5.1]. By Proposition 2.14, the proper intersection product on the right is symmetric with respect to the Cartier divisors.

Let 𝒴 , 𝒴 ′ be varieties over K^{∘} and let φ : 𝕋 ×_{K^{∘}} 𝒴 → 𝕋 ×_{K^{∘}} 𝒴 ′ be a flat morphism over 𝕋. The point t ∈ T^{∘}(K) corresponds to an integral point of 𝕋 inducing a flat morphism φ_{t} : 𝒴 → 𝒴 ′ by base change from φ. We recall from 2.5 that we have also introduced the pull-back with respect to flat morphisms. The following result shows some functoriality with the above pull-backs. We will use the canonical projection p_{2} : 𝕋 ×_{K^{∘}} 𝒴 ′ → 𝒴 ′.

Proposition 5.5. Under the hypothesis above, let 𝒵′ be a cycle of 𝒴 ′. Then the cycle φ^{∗}(p_{2}^{∗}(𝒵′)) satisfies the flatness condition from 5.3 and we have i_{t}^{∗}(φ^{∗}(p_{2}^{∗}(𝒵′))) = φ_{t}^{∗}(𝒵′).

Proof. The cycle p_{2}^{∗}(𝒵) satisfies the flatness condition. Using that φ is a flat morphism over 𝕋, we deduce that φ^{∗}(p_{2}^{∗}(𝒵′)) also fulfills the flatness condition. Since the pull-backs are defined for horizontal and vertical parts in terms of the corresponding operations for varieties over fields, the claim follows from [9, Proposition 6.5], where the functoriality for pull-backs of cycles of varieties has been proved.

Lemma 5.6. Let 𝒴 be a variety over K^{∘} and let t ∈ T^{∘}(K). Suppose that g is a rational function on 𝕋 ×_{K^{∘}} 𝒴 such that every irreducible component of the support of the restriction of div(g) to the generic fibre is flat over T. Then g(t, ·) is a rational function on 𝒴 and we have i_{t}^{∗}(cyc(div(g))) = cyc(div(g(t, ·))).

Proof. The flatness assumption yields the first claim immediately. Note that the vertical components of cyc(div(g)) are automatically flat over 𝕋_{s} and hence we get a well-defined cycle i_{t}^{∗}(cyc(div(g))) on 𝒴 . The second claim follows easily from the fact that we may write i_{t}^{∗} as an n-fold proper intersection product with Cartier divisors (see 5.4) and from Proposition 2.14.

Proposition 5.7. Let 𝒴 be a variety over K^{∘}. Then pull-back with respect to the canonical projection p_{2} : 𝕋 ×_{K^{∘}} 𝒴 → 𝒴 induces an isomorphism p_{2}^{∗} : CH^{1}(𝒴 ) → CH^{1}(𝕋 ×_{K^{∘}} 𝒴 ).

Proof. By 2.16, p_{2}^{∗} is compatible with rational equivalence and hence it is welldefined on the Chow groups. Surjectivity follows from Proposition 5.2. Suppose that 𝒟 is a cycle of codimension 1 on 𝒴 such that p_{2}^{∗}(𝒟) is rationally equivalent to 0 on 𝕋×_{K^{∘}} 𝒴 . Using Lemma 5.6 for the unit element e in T^{∘}(K), we deduce that 𝒟 is rationally equivalent to 0. This proves injectivity.

We have a similar statement for Picard group as pointed out by Qing Liu and C. Pépin.

Proposition 5.8. Let 𝒴 be a normal variety over K^{∘}. Then pull-back with respect to p_{2} induces an isomorphism Pic(𝒴 ) → Pic(𝕋 ×_{K^{∘}} 𝒴 ).

Proof. This statement was proved in [14, Remark 9.6] based on an argument of Qing Liu and C. Pépin.

Now let 𝒴 be a normal 𝕋-toric variety over K^{∘} and let 𝒟 be a cycle of codimension 1 in 𝒴 . Note that t ∈ T^{∘}(K) acts on 𝒴 and we denote by 𝒟^{t} the pull-back of 𝒟 with respect to this flat morphism.

Proposition 5.9. Under the hypothesis above, 𝒟^{t} is rationally equivalent to 𝒟.

Proof. Let σ : 𝕋 ×_{K^{∘}} 𝒴 → 𝒴 be the torus action on 𝒴 . It follows from Propositions 5.2 that σ^{∗}(𝒟) is rationally equivalent to p_{2}^{∗}(𝒟′) for a cycle 𝒟′ of codimension 1 in 𝒴 . Then Lemma 5.6 and Proposition 5.5, applied for the unit element e, show that 𝒟 is rationally equivalent to 𝒟′. By 2.16, we conclude that σ^{∗}(𝒟) is rationally equivalent to p_{2}^{∗}(𝒟). If we apply Proposition 5.5 again, but now in t instead of e, we get the claim.

Lemma 5.10. Let 𝒰_{0} be a non-empty open subset of the 𝕋-toric variety 𝒴 over K^{∘} and let 𝒰 := ⋃_{t∈T^{∘}(K)} t𝒰_{0}. Then 𝒰 is the smallest 𝕋-invariant (open) subset containing 𝒰_{0}.

Proof. Consider the subset S of 𝕋 such that translation with its elements leaves 𝒰 invariant. The subset S ∩𝕋_{s} is equal to the stabilizer of 𝒴_{s} \𝒰_{s} and hence it is an algebraic subgroup of 𝕋_{s}. By construction, it contains 𝕋_{s}(K̃) and hence it is equal to 𝕋_{s}. We use the same argument for the points of S contained in the generic fibre T = 𝕋_{η}. Again, S ∩ T is an algebraic subgroup containing T^{∘}(K). Since T^{∘} is an n-dimensional affinoid torus, we conclude that T^{∘}(K) is Zariski dense in T and hence the algebraic subgroup is the torus T over K. We conclude that 𝒰 is 𝕋-invariant. This proves the claim immediately.

5.11. Since the torus 𝕋_{s} acts continuously on the discrete set of the generic points of 𝒴_{s}, every such generic point is fixed under the action. We conclude that every irreducible component of 𝒴_{s} is invariant under the 𝕋-action. This means that the special fibres of 𝒰 and 𝒰_{0} have the same generic points. We have seen in Proposition 5.1 that 𝒰 \ 𝒰_{0} is a union of irreducible components of codimension 1 and hence every such irreducible component is horizontal. Let 𝒟 be the horizontal cycle on 𝒰 given by the formal sum of these irreducible components.

Proposition 5.12. Under the hypothesis above, there is a unique Cartier divisor D on 𝒰 such that 𝒟 = cyc(D). Moreover, this Cartier divisor is effective.

Proof. For t ∈ T^{∘}(K), Proposition 5.9 yields a non-zero rational function f_{t} on 𝒰 such that 𝒟 − 𝒟^{t} = cyc(div(f_{t})). Since 𝒰 \ t^{−1}𝒰_{0} is equal to the support of 𝒟^{t}, we deduce that the restriction of 𝒟 to t^{−1}𝒰_{0} is the Weil divisor given by the rational function f_{t} on t^{−1}𝒰_{0}. By Corollary 2.12, the Cartier divisor on a normal variety is uniquely determined by its associated Weil divisor. This yields immediately that {(t^{−1}𝒰_{0}, f_{t}) | t ∈ T^{∘}(K)} is a Cartier divisor on 𝒰 with associated Weil divisor 𝒟 and uniqueness follows as well. By Corollary 2.12, the Cartier divisor D is effective.

Proposition 5.13. Let σ : 𝕋 ×_{K^{∘}} 𝒴 → 𝒴 be the torus action of the normal 𝕋-toric variety 𝒴 over K^{∘} and let D be the Cartier divisor from Proposition 5.12. Then σ^{∗}(D) is linearly equivalent to p_{2}^{∗}(D).

Proof. The unit element e in T^{∘}(K) induces the section i_{e} of σ and p_{2}. Then the claim follows from Proposition 5.8. Another way to deduce the claim is to use the corresponding statement for cycles of codimension 1 (see Proposition 5.7) together with Corollary 2.12.

Corollary 5.14. Let D be the Cartier divisor from Proposition 5.12 and let D^{t} be its pull-back with respect to translation by t ∈ T^{∘}(K). Then the invertible sheaves 𝒪(D^{t}) and 𝒪(D) on 𝒰 are isomorphic. Moreover, 𝒪(D) is generated by global sections.

Proof. The first claim follows from Proposition 5.13 by applying i_{t}^{∗}. By Proposition 5.12, s_{D^{t}} is a global section with support 𝒰 \t^{−1}𝒰_{0} and hence the second claim follows from the first.

6 Linearization and immersion into projective space

Let 𝕋 be the split torus of rank n over K^{∘} and let 𝒴 be a normal 𝕋-toric variety over K^{∘}. We denote by μ : 𝕋 ×_{K^{∘}} 𝕋 → 𝕋 the multiplication map, by σ : 𝕋 ×_{K^{∘}} 𝒴 → 𝒴 the group action and by p_{2} : 𝕋 ×_{K^{∘}} 𝒴 → 𝒴 the second projection. As in the previous section, we consider a non-empty affine open subset 𝒰_{0} of 𝒴 and the smallest 𝕋-invariant open subset 𝒰 of 𝒴 containing 𝒰_{0}. In Proposition 5.12, we have constructed an effective Cartier divisor D on 𝒰 with supp(D) = 𝒰 \ 𝒰_{0} such that cyc(D) is a horizontal cycle with all multiplicities equal to 1. In this section, we will see that 𝒪(D) has a 𝕋linearization (Proposition 6.3) and is ample (Proposition 6.7) leading in 6.8 to a 𝕋-equivariant immersion into a projective space with linear 𝕋-action. This will be summarized in Proposition 6.9 which reduces the proof of Sumihiro’s theorem to an easier projective variant shown in the next section.

Definition 6.1. First, we recall the definition of a 𝕋-linearization of a line bundle L on a toric variety (see [20] for details). Geometrically, a 𝕋-linearization is a lift of the torus action on 𝒴 to an action on L such that the zero section is 𝕋-invariant. In terms of the underlying invertible sheaf ℒ, a 𝕋-linearization is an isomorphism

ϕ : σ^{∗}ℒ → p_{2}^{∗}ℒ ,

of sheaves on 𝕋 ×_{K^{∘}} 𝒴 satisfying the cocycle condition

p_{23}^{∗}ϕ ∘ (id_{𝕋} × σ)^{∗}ϕ = (μ × id_{𝒴} )^{∗}ϕ,  (7)

where p_{23} : 𝕋×_{K^{∘}} 𝕋×_{K^{∘}} 𝒴 → 𝕋×_{K^{∘}} 𝒴 is the projection to the last two factors.

We need the following application of a result of Rosenlicht.

Lemma 6.2. For every f ∈ 𝒪(𝕋 ×_{K^{∘}} 𝒰 )^{×}, there is a character χ on T and a g ∈ 𝒪(𝒰 )^{×} such that f = χ · g.

Proof. Note that 𝒪(𝕋)^{×} is the set of characters on T multiplied by units in K^{∘}. Then the claim follows from [23, Theorem 2], where it is proved in the case of fields.

Proposition 6.3. The invertible sheaf 𝒪(D) has a 𝕋-linearization.

Proof. By Proposition 5.13, we have an isomorphism

ϕ : σ^{∗}ℒ → p_{2}^{∗}ℒ

for the invertible sheaf ℒ := 𝒪(D). Both sides of (7) are isomorphisms between the same invertible sheaves on 𝕋 ×_{K^{∘}} 𝕋 ×_{K^{∘}} 𝒰 and hence there is a unique f ∈ 𝒪(𝕋 ×_{K^{∘}} 𝕋 ×_{K^{∘}} 𝒰 )^{×} such that that the left hand side is obtained by multiplying the right hand side with f. We may choose ϕ such that we have the canonical isomorphism over {e} × 𝒰 and hence we get f(e, ·, ·) = 1 and f(·, e, ·) = 1. By Lemma 6.2, there are characters χ_{1}, χ_{2} on T and g ∈ 𝒪(𝒰 )^{×} such that f(t_{1}, t_{2}, u) = χ_{1}(t_{1})χ_{2}(t_{2})g(u) for all t_{1}, t_{2} ∈ T (\overline{K}) and u ∈ 𝒰 (\overline{K}). Since f(e, e, u) = 1, we get g = 1. Therefore

f(t_{1}, t_{2}, u) = χ_{1}(t_{1})χ_{2}(t_{2}) = f(t_{1}, e, u)f(e, t_{2}, u) = 1.

By density of the \overline{K}-rational points, we get f = 1 and (7) holds.

6.4. The 𝕋-linearization on ℒ = 𝒪(D) induces a dual action of 𝕋 on the space H^{0}(𝒰 , L) of global sections, given by the composition σ̂ of the canonical K^{∘}-linear maps

H^{0}(𝒰 , ℒ) → H^{0}(𝕋 ×_{K^{∘}} 𝒰 , σ^{∗}ℒ) → → H^{0}(𝕋 ×_{K^{∘}} 𝒰 , p_{2}^{∗}ℒ) → H^{0}(𝕋, 𝒪_{𝕋}) ⊗_{K^{∘}} H^{0}(𝒰 , ℒ),

where the last isomorphism comes from the Künneth formula (see [16]). We refer to [20, Chapter 1, Definition 1.2] for the definition of a dual action. This was written for vector spaces over a base field, but the same definition applies in case of a free K^{∘}-module. Since V := H^{0}(𝒰 , ℒ) is a torsion free K^{∘}-module, V is indeed free (see [14, Lemma 4.2] for a proof). A dual action means that the torus 𝕋 acts linearly on the possibly infinite dimensional projective space ℙ(V) = Proj(K^{∘}[V]). The dual action σ̂ induces an action of t ∈ T^{∘}(K) on V which we denote by s ↦ t · s. For s ∈ V = H^{0}(𝒰 , ℒ), the action is geometrically given by (t · s)(u) = t^{−1}(s(tu)), u ∈ 𝒰 , where t^{−1} operates on the underlying line bundle using the linearization.

Lemma 6.5. Let x_{1}, …, x_{k} be affine coordinates of 𝒰_{0} considered as rational functions on 𝒰 . Then there exists ℓ ∈ ℤ_{+} such that for every i ∈ {1, …, k}, the meromorphic section s_{i} := x_{i}s_{ℓD} of 𝒪(ℓD) is in fact a global section. Here, s_{ℓD} denotes the canonical global section of 𝒪(ℓD).

Proof. Using the theory of divisors from §2, we get the identity

cyc(div(x_{i})) = ∑_{j} m_{ij}Z_{j} + 𝒱

of cycles on 𝒰 , where Z_{j} are the irreducible components of 𝒰 \𝒰_{0} and where 𝒱 is an effective cycle of codimension 1 in 𝒰 which meets 𝒰_{0}. By construction (see Proposition 5.12), we have cyc(D) = 𝒟 = ∑_{j} Z_{j}. For ℓ := − min_{ij}{m_{ij}, 0}, we get

cyc(div(x_{i})) + ℓ𝒟 = ∑_{j} m_{ij}Z_{j} + ℓ𝒟 + 𝒱 = ∑_{j}(m_{ij} + ℓ)Z_{j} + 𝒱 ≥ 0.

Therefore the Weil divisor div(x_{i}) + ℓD is effective. By Corollary 2.12, we conclude that x_{i}s_{ℓD} is a global section of 𝒪(ℓD).

6.6. By Proposition 2.2, 𝒴 is a noetherian topological space and therefore 𝒰 is quasicompact. Using Lemma 5.10, there is a finite subset S of T^{∘}(K) such that 𝒰 = ⋃_{t∈S} t^{−1}𝒰_{0}. We have seen in Lemma 6.5 that the affine coordinates x_{1}, …, x_{k} of 𝒰_{0} induce global sections s_{1}, …, s_{k} of 𝒪(ℓD). Then the dual action from 6.4 gives global sections t · s_{1}, …, t · s_{k} of 𝒪(ℓD) induced by affine coordinates of t^{−1}𝒰_{0}. We conclude that (t · s_{j})_{t∈S,j=1,...,k} generate 𝒪(ℓD). By construction, the global section t · s_{D} has support 𝒰 \ t^{−1}𝒰_{0}. We get a morphism

ψ : 𝒰 → ℙ_{K^{∘}}^{R′}, u ↦ (⋯: t · s_{1}(u) : ⋯: t · s_{k}(u) : t · s_{D}^{ℓ}(u) : ⋯)_{t∈S}

with R′ := |S|(k + 1) − 1. Note that this map is well defined because 𝒪(ℓD) is generated by these global sections, and we have ψ^{∗}(𝒪_{ℙ_{K^{∘}}^{R′}} (1)) ≃ 𝒪(ℓD).

Proposition 6.7. The morphism ψ is an immersion and hence ℒ is ample.

Proof. For t ∈ S, the support of the Cartier divisor div(t · s_{D}) = D^{t} is equal to 𝒰 \t^{−1}𝒰_{0}. Let y_{j} be the coordinate of ℙ_{K^{∘}}^{R′} corresponding to t·s_{D}^{ℓ} with respect to the morphism ψ. Then we get ψ^{−1}{y_{j} ≠ 0} = t^{−1}𝒰_{0}. Since t · x_{1}, …, t · x_{k} are affine coordinates on t^{−1}𝒰_{0}, we conclude easily that ψ restricts to a closed immersion of t^{−1}𝒰_{0} into the open subvariety {y_{j} ≠ 0} of ℙ_{K^{∘}}^{R′}. Since these open subvarieties form an open covering of ℙ_{K^{∘}}^{R′}, we may use [11, Corollaire 4.2.4] to conclude that the morphism ψ is an immersion and hence 𝒪(D) is ample.

6.8. Let V_{ℓ0} be the submodule of V_{ℓ} := H^{0}(𝒰 , ℒ^{ℓ}) which is generated by the global sections (t·s_{j})_{t∈S,j=1,...,k} and (t·s_{D}^{ℓ})_{t∈S} used in the definition of ψ in 6.6. Since 𝒪(ℓD) has a 𝕋-linearization, we get a dual action σ̂ of 𝕋 on V_{ℓ} similarly to 6.4. A K^{∘}-submodule W of V_{ℓ} is called invariant under the dual action of 𝕋 if σ̂(W) ⊂ A ⊗_{K^{∘}} W for A := H^{0}(𝕋, 𝒪_{𝕋}) = K^{∘}[M]. The lemma on p. 25 of [20] generalizes in a straightforward manner to our setting and hence there is a finitely generated submodule W of V_{ℓ} which is invariant under the dual action of 𝕋 and contains V_{ℓ0}. Since W is K^{∘}-torsion free, we conclude that W is a free K^{∘}-module of finite rank R + 1. We get a morphism i : 𝒰 → ℙ(W) with i^{∗}(𝒪_{ℙ(W)}(1)) ≅ 𝒪(ℓD). The dual action of 𝕋 on W induces a linear action of 𝕋 on the projective space ℙ(W).

By construction, i is 𝕋-equivariant. Since i factorizes through ψ, we deduce from Proposition 6.7 that i is an immersion.

Recall that 𝕋 = Spec(K^{∘}[M]) is the split torus of rank n. We summarize our findings:

Proposition 6.9. Let 𝒰_{0} be a non-empty affine open subset of the normal 𝕋-toric variety 𝒴 over K^{∘} and let 𝒰 be the smallest 𝕋-invariant open subset of 𝒴 containing 𝒰_{0}. Then there is a 𝕋-equivariant open immersion of 𝒰 into a projective 𝕋-toric variety 𝒴_{A,a} given by A ∈ M^{R+1} and height function a as in 3.5.

Proof. Let i : 𝒰 → ℙ(W) be the 𝕋-equivariant immersion from 6.8. Then the closure 𝒴 of i(𝒰 ) in ℙ(W) is a projective 𝕋-toric variety over K^{∘} on which 𝕋-acts linearly. We choose a K-rational point y in the open dense orbit of i(𝒰 ). By [14, Proposition 9.8], there are suitable coordinates on ℙ(W) and A ∈ M^{R+1} such that 𝒴 = 𝒴_{A,a} for the height function a of y defined in 3.5.

7 Proof of Sumihiro’s theorem

In this section, we will finally prove Sumihiro’s theorem for normal toric varieties over K^{∘} as announced in Theorem 2. Sumihiro’s theorem is wrong for arbitrary non-normal toric varieties even over a field (see [6, Example 3.A.1] for a projective counterexample). However, we will show first in this section that Sumihiro’s theorem holds for open invariant subsets of projective toric varieties over K^{∘} with a linear torus action. Note that such projective toric varieties are not necessarily normal. As a consequence of Proposition 6.9, we will obtain Sumihiro’s theorem for normal toric varieties over K^{∘}. At the end, Theorem 3 will easily follow from Sumihiro’s theorem similarly as in the classical case of normal toric varieties over a field. In this section, (K, v) will be a valued field with value group Γ = v(K^{×}) ⊂ ℝ. Moreover, 𝕋 = Spec(K^{∘}[M]) is a split torus over the valuation ring K^{∘} of rank n and we consider a projective 𝕋-toric variety over K^{∘} with a linear 𝕋-action. By [14, Proposition 9.8] the latter is a toric subvariety 𝒴_{A,a} of ℙ_{K^{∘}}^{R} as in 3.5 for A ∈ M^{R+1}, height function a and suitable projective coordinates x_{0}, …, x_{R}.

7.1. We fix a point z ∈ 𝒴_{A,a} and a closed 𝕋-invariant subset Y of 𝒴_{A,a} with z ∉ Y. Since Y is a closed subset of the ambient projective space ℙ_{K^{∘}}^{R}, there is a k ∈ ℤ_{+} and s_{0} ∈ H^{0}(ℙ_{K^{∘}}^{R}, 𝒪(k)) such that s_{0}|_{Y} = 0 and s_{0}(z) ≠ 0. Obviously, V := H^{0}(ℙ_{K^{∘}}^{R}, 𝒪(k)) is a free K^{∘}-module of finite rank. The linear 𝕋-action on ℙ_{K^{∘}}^{R} induces a linear representation of 𝕋 on V, i.e. a homomorphism S : 𝕋 → GL(V) of group schemes over K^{∘}. We say that s ∈ V is semi-invariant if there is u ∈ M such that S_{t}(s) = χ^{u}(t)s for every t ∈ 𝕋 and for the character χ^{u} of 𝕋 associated to u. In the following, the K^{∘}-submodule

W := {s ∈ H^{0}(ℙ_{K^{∘}}^{R}, 𝒪(k)) | ∃λ ∈ K^{∘}\{0} s. t. λs|_{Y} = 0}

of V will be of interest. Note that W is equal to the set of global sections s of 𝒪(k) which vanish on the generic fibre Y_{η}. Since Y is 𝕋-invariant, it is clear that W is invariant under the 𝕋-action. The multiplicative torus T = 𝕋_{K} is split over K and hence the vector space W_{K} has a simultaneous eigenbasis for the T -action. This is well-known in the theory of toric varieties over a field and follows from [2, Proposition III.8.2]. Note that this K-basis is semi-invariant. We will show in the next lemma that such a basis exists as a basis of W over K^{∘}.

Lemma 7.2. W is a free K^{∘}-module of finite rank which has a semi-invariant basis.

Proof. A valuation ring is a Prüfer domain. Since W is a saturated K^{∘}- submodule of the free module V of finite rank, we conclude that W is free of finite rank r (see [4, ch. VI, §4, Exercise 16]). We have seen above that W_{K} has a simultaneous eigenbasis w_{1}, …, w_{r} for the T -action. For j = 1, …, r, we have S_{t}(w_{j}) = χ^{u_{j}}(t) · w_{j} for all t ∈ T (K) and some u_{j} ∈ M. Let E_{u_{j}} be the corresponding eigenspace. Then W_{u_{j}} := E_{u_{j}} ∩ V is a saturated K^{∘}-submodule of W. The same argument as above shows that W_{u_{j}} is a free K^{∘}-module of finite rank. We may choose the simultaneous eigenbasis w_{1}, …, w_{r} above in such a way that a suitable subset is a K^{∘}-basis of W_{u_{j}} for every j = 1, …, r. Note that every w_{j} is semi-invariant. For t in the subgroup U := 𝕋(K^{∘}) = T^{∘}(K) of T (K), we have S_{t} ∈ GL(V, K^{∘}) and hence the eigenvalues χ^{u_{j}}(t) have valuation 0. Using reduction modulo the maximal ideal K^{∘∘} of K^{∘}, the U-action becomes a 𝕋K̃-operation on W̃ := W ⊗_{K^{∘}} K̃. We note that the reduction of a K^{∘}-basis in W_{u_{j}} is linearly independent in W̃. Using that eigenvectors for distinguished eigenvalues are linearly independent, we conclude that the reduction of w_{1}, …, w_{r} is a a simultaneous eigenbasis for the 𝕋K̃-action on W̃. By Nakayama’s Lemma, it follows that w_{1}, …, w_{r} is a K^{∘}-basis for W.

We can now prove the following quasi-projective version of Sumihiro’s theorem.

Proposition 7.3. Let 𝒰 be a 𝕋-invariant open subset of 𝒴_{A,a}. Then every point of 𝒰 has a 𝕋-invariant open affine neighbourhood in 𝒰 .

Proof. Let z ∈ 𝒰 and let Y := 𝒴 \ 𝒰 . Since Y is 𝕋-invariant, we are in the setting of 7.1 and we will use the notation from there. In particular, we have s_{0} ∈ H^{0}(ℙ_{K^{∘}}^{R}, 𝒪(k)) such that s_{0}|_{Y} = 0 and s_{0}(z) ≠ 0. Using Lemma 7.2, we conclude that there is a semi-invariant s_{1} ∈ H^{0}(ℙ_{K^{∘}}^{R}, 𝒪(k)) with s_{1}(z) ≠ 0 and λs_{1}|_{Y} = 0 for some λ ∈ K^{∘} \ {0}. To construct the affine invariant neighborhood of z, we assume first that z is contained in the generic fibre of 𝒰 over K^{∘}. Then 𝒰_{1} := {x ∈ 𝒴_{A,a} | λs_{1}(x) ≠ 0} is an affine open subset of 𝒰 that contains z. Since s_{1} is semi-invariant, it follows that 𝒰_{1} is 𝕋-invariant proving the claim.

Now we suppose that z is contained in the special fibre 𝒰_{s}. Let ζ be a generic point of an irreducible component Z of Y_{s}. Since Y is 𝕋-invariant, ζ is the generic point of an orbit whose closure Z does not contain z. Using the orbit– face correspondence from Proposition 3.6, there is a projective coordinate x_{i(ζ)} such that x_{i(ζ)}(Z) = 0 but x_{i(ζ)}(z) ≠ 0. By definition of 𝒴_{A,a}, we may view x_{i(ζ)} as a semi-invariant global section of 𝒪(1) on ℙ_{K^{∘}}^{R}. Letting ζ varying over the generic points of the irreducible components of Y_{s}, we get a semi-invariant global section s := s_{1} · ∏_{ζ} s_{i(ζ)} of a suitable tensor power of 𝒪(1) on ℙ_{K^{∘}}^{R} with s(z) ≠ 0 and s|_{Y} = 0. Then 𝒰_{1} := {x ∈ 𝒴_{A,a} | s(x) ≠ 0} is a 𝕋-invariant affine open neighbourhood of z in 𝒰 .

Proof of Theorem 2. We are now ready to prove Sumihiro’s theorem for a normal 𝕋-toric variety 𝒴 over K^{∘}. Every point z ∈ 𝒴 has an affine open neighbourhood 𝒰_{0}. Let 𝒰 be the smallest 𝕋-invariant open subset of 𝒴 containing 𝒰_{0}. By Proposition 6.9, there is an equivariant open immersion i : 𝒰 → 𝒴_{A,a} for suitable A ∈ M^{R+1} and height function a. By Proposition 7.3, there is a 𝕋-invariant open neighbourhood 𝒰_{1} of z in i(𝒰 ). We conclude that i^{−1}(𝒰_{1}) is an affine 𝕋-invariant open neighbourhood of z in 𝒰 proving Sumihiro’s theorem.

Finally in order to complete the picture which gives rise to the interplay between toric geometry and convex geometry, we prove Theorem 3 which give us a bijective correspondence between normal 𝕋-toric varieties and Γ-admissible fans.

Proof of Theorem 3. We assume first that v is not a discrete valuation. For simplicity, we fix torus coordinates on the split torus 𝕋 of rank n. Let 𝒴 be a normal 𝕋-toric variety. By Theorem 2, 𝒴 has an open covering {𝒱_{i}}_{i∈I} by affine 𝕋-varieties 𝒱_{i}. By Theorem 1, we have 𝒱_{i} ≅ 𝒱_{σ_{i}} for a Γ-admissible cone σ_{i} in ℝ^{n} ×ℝ_{+} for which the vertices of σ_{i} ∩(ℝ^{n} ×{1}) are contained in Γ^{n} ×{1}. Since 𝒴 is separated, 𝒱_{ij} := 𝒱_{i}∩𝒱_{j} is affine for every i, j ∈ I. We conclude that 𝒱_{i} ∩ 𝒱_{j} is an affine normal 𝕋-toric variety and hence Theorem 1 again shows 𝒱_{ij} ≅ 𝒱_{σ_{ij}} for a Γ-admissible cone σ_{ij} in ℝ^{n} × ℝ_{+}. Applying the orbit–face correspondence from [14, Proposition 8.8] to the open immersions 𝒱_{ij} → 𝒱_{i} and 𝒱_{ij} → 𝒱_{j}, it follows that σ_{ij} is a closed face of σ_{i} and σ_{j}. Moreover, the same argument shows that σ_{ij} = σ_{i} ∩ σ_{j} and hence the closed faces of all σ_{i} form a Γ-admissible fan Σ in ℝ^{n} × ℝ_{+} with 𝒴_{Σ} ≅ 𝒴 . From Theorem 1, we get now easily the desired bijection. If v is a discrete valuation, then the same argument works if we omit the additional condition on the vertices of the cones.

References

[1] Vladimir G. Berkovich. Spectral theory and analytic geometry over non-Archimedean fields, volume 33 of Mathematical Surveys and Monographs. American Mathematical Society, Providence, RI, 1990.

[2] Armand Borel. Linear algebraic groups, volume 126 of Graduate Texts in Mathematics. Springer-Verlag, New York, second edition, 1991.

[3] S. Bosch, U. Güntzer, and R. Remmert. Non-Archimedean analysis, volume 261 of Grundlehren der Mathematischen Wissenschaften [Fundamental Principles of Mathematical Sciences]. Springer-Verlag, Berlin, 1984. A systematic approach to rigid analytic geometry.

[4] Nicolas Bourbaki. Commutative algebra. Chapters 1–7. Elements of Mathematics (Berlin). Springer-Verlag, Berlin, 1989. Translated from the French, Reprint of the 1972 edition.

[5] J.I. Burgos Gil, P. Philippon, and M. Sombra. Arithmetic Geometry of Toric Varieties. Metrics, Measures and Heights. Astérisque 360, 2014.

[6] David A. Cox, John B. Little, and Henry K. Schenck. Toric varieties, volume 124 of Graduate Studies in Mathematics. American Mathematical Society, Providence, RI, 2011.

[7] Günter Ewald. Combinatorial convexity and algebraic geometry, volume 168 of Graduate Texts in Mathematics. Springer-Verlag, New York, 1996.

[8] William Fulton. Introduction to toric varieties, volume 131 of Annals of Mathematics Studies. Princeton University Press, Princeton, NJ, 1993. The William H. Roever Lectures in Geometry.

[9] William Fulton. Intersection theory, volume 2 of Ergebnisse der Mathematik und ihrer Grenzgebiete. 3. Folge. A Series of Modern Surveys in Mathematics [Results in Mathematics and Related Areas. 3rd Series. A Series of Modern Surveys in Mathematics]. Springer-Verlag, Berlin, second edition, 1998.

[10] I. M. Gel′fand, M. M. Kapranov, and A. V. Zelevinsky. Discriminants, resultants, and multidimensional determinants. Mathematics: Theory & Applications. Birkhäuser Boston Inc., Boston, MA, 1994.

[11] A. Grothendieck. Éléments de géométrie algébrique. I. Le langage des schémas. Inst. Hautes Études Sci. Publ. Math., (4):228, 1960.

[12] Walter Gubler. Local heights of subvarieties over non-Archimedean fields. J. Reine Angew. Math., 498:61–113, 1998.

[13] Walter Gubler. Local and canonical heights of subvarieties. Ann. Sc. Norm. Super. Pisa Cl. Sci. (5), 2(4):711–760, 2003.

[14] Walter Gubler. A guide to tropicalizations. In Algebraic and combinatorial aspects of tropical geometry, volume 589 of Contemp. Math., pages 125– 189. Amer. Math. Soc., Providence, RI, 2013.

[15] G. Kempf, Finn Faye Knudsen, D. Mumford, and B. Saint-Donat. Toroidal embeddings. I. Lecture Notes in Mathematics, Vol. 339. Springer-Verlag, Berlin, 1973.

[16] George R. Kempf. Some elementary proofs of basic theorems in the cohomology of quasicoherent sheaves. Rocky Mountain J. Math., 10(3):637–645, 1980.

[17] Hagen Knaf. Divisors on varieties over valuation domains. Israel J. Math., 119:349–377, 2000.

[18] Diane Maclagan and Bernd Sturmfels. Introduction to Tropical Geometry, volume 161 of Graduate Studies in Mathematics. American Mathematical Society, Providence, RI, 2015.

[19] Grigory Mikhalkin. Tropical geometry and its applications. In International Congress of Mathematicians. Vol. II, pages 827–852. Eur. Math. Soc., Zürich, 2006.

[20] D. Mumford, J. Fogarty, and F. Kirwan. Geometric invariant theory, volume 34 of Ergebnisse der Mathematik und ihrer Grenzgebiete (2) [Results in Mathematics and Related Areas (2)]. Springer-Verlag, Berlin, third edition, 1994.

[21] Tadao Oda. Convex bodies and algebraic geometry, volume 15 of Ergebnisse der Mathematik und ihrer Grenzgebiete (3) [Results in Mathematics and Related Areas (3)]. Springer-Verlag, Berlin, 1988. An introduction to the theory of toric varieties, Translated from the Japanese.

[22] Fred Rohrer. The geometry of toric schemes. J. Pure Appl. Algebra, 217(4):700–708, 2013.

[23] Maxwell Rosenlicht. Toroidal algebraic groups. Proc. Amer. Math. Soc., 12:984–988, 1961.

[24] A. L. Smirnov. Torus schemes over a discrete valuation ring. Algebra i Analiz, 8(4):161–172, 1996.

[25] Hideyasu Sumihiro. Equivariant completion. J. Math. Kyoto Univ., 14:1– 28, 1974.

[26] Peter Ullrich. The direct image theorem in formal and rigid geometry. Math. Ann., 301(1):69–104, 1995.

Walter Gubler Fakultät für Mathematik Universität Regensburg Universitätsstrasse 31 D-93040 Regensburg walter.gubler@mathematik.uniregensburg.de Alejandro Soto Department of Mathematics KU Leuven Celestijnenlaan 200B B-3001 Heverlee alejandro.soto@wis.kuleuven.be